OpenStudy (anonymous):

Use the substitution method to find the solution(s) to the system of equations. Show all work for full credit. y = x^2 y = x + 2 Use the substitution method to find the solution(s) to the system of equations. Show all work for full credit. y = x^2 y = x + 2 @Mathematics

OpenStudy (anonymous):

-x - 1 = x² - 1 Then, 0 = x² - 1 + x + 1 0 = x² + x 0 = x(x + 1) x = {0,-1} Next, by substituting those values for either equation: y = -0 - 1 = -1 y = -(-1) - 1 = 0 Hence, the solutions are (0,-1) and (-1,0)

OpenStudy (anonymous):

\[x^2 = x+2\]\[x^2 - x - 2 = 0\]\[(x -2)(x+1) =0\]\[x = 2, -1\]\[substitute => (2,4), (-1,1)\]